# Intermittent localization and fast spatial learning by non-Markov random walks with decaying memory

### v1

@@ VIII Conclusions
Previous work has indicated the existence of localized states for certain random walks with preferential resetting to sites visited in the past [20, 21, 22]. A simple model considers an infinite $1d$ lattice, with a single impurity site representing a resource patch with a larger attractiveness than the other sites. Due to the reinforced dynamics, the walker does not diffuse away at late times but tends to an exponential stationary distribution peaked at the impurity. This emergent property can be associated to a phenomenon of spatial learning, mimicking the activity of a foraging animal that would settle down thanks to memory around a zone where resources are abundant in a scarce environment. In the present work, we have studied how such a localization is affected by imperfect memory, when the walker slowly forgets about the steps performed further into in the past. This assumption is more realistic for animals and humans and also gives rise to new phenomena.
@@
As a first result, we have observed that if memory decays slowly enough, such that the effective number of remembered time-steps still grows unbounded with time (albeit sub-linearly), the walker keeps localizing and its stationary density is actually the same as in the perfect memory case. Secondly, according to a de-correlation approximation, these localized solutions turn out to be linearly stable. The relaxation of small perturbations depends on the memory kernel and it is not exponential with time: the leading mode relaxes to zero as an inverse power-law of time $t^{-|\nu|}$ with a non-trivial exponent which depends on the parameters of the model. A power-law is also thought to describe the relaxation of the classic vertex reinforced random walk, in connection with urn models [7].
@@
Thirdly, a memory loss of the form of $1/\tau$ represents a boundary case which displays interesting peculiarities: it is the memory that decays the fastest and still allows the same localization as the perfect memory case. In other words, this memory is the “cheapest”, as the effective number of remembered steps up to time $t$ only grows as $\ln t$. In addition, somehow counter-intuitively, a walker using the $1/\tau$-kernel exhibits the fastest relaxation toward the localized distribution, see Eq. (58). Hence, such a walker learns faster than a walker with better memory. The speed of learning measured through the exponent $|\nu|$ [obtained from Eq. (53)] increases as $\beta$, the exponent that describes the decay of memory, varies from 0 to 11. This result resonates with empirical observations in the neurosciences indicating that forgetting is essential to the brain and can actually be beneficial to learning processes [37, 38, 39]. Memory kernels in $1/\tau$ have been inferred from bison foraging data [32] and can emerge in mean-field models of interacting synapses subject to Hebbian connections and competition [46].
@@
As a fourth result, spatial learning starts to be disrupted when the memory kernel decays exponentially, or faster than $1/\tau$, i.e., when the effective number of remembered steps saturates to a constant value at late times. A new regime of intermittent localization appears, which consists of random time intervals of localization separated by random diffusive excursions that wander away from the impurity. Numerically, the duration of these localized periods is exponentially distributed whereas that of the diffusive excursions follows a power-law distribution with infinite first moment. Consequently, over a very long observation window, the walker will tend to be found mostly in the non-localized state. Nevertheless, as the parameter of memory use, $q$, increases beyond a crossover value, the mean duration of the localized periods becomes extremely large. Remarkably, during a localization period, the distribution of the walker’s position is practically undistinguishable from the localized profile with perfect memory, even when the localization periods are relatively short. Unfortunately, in the intermittent localization regime, the de-correlation approximation breaks down and is not able to predict the loss of stability (or the absence) of stationary solutions.
@@
Intermittency in low-dimensional deterministic dynamical systems is well-known to occur through a saddle-node bifurcation, beyond which an attractor fixed point no longer exists, allowing intermittent bursts of chaos [47]. Just at the bifurcation, the relaxation toward the marginal fixed point follows a power-law form in $1/t$, which is similar, although not strictly identical, to Eq. (58) for our boundary case. The $1/t$ relaxation at the edge of intermittency in deterministic systems corresponds to a critical slowing down, which is generic at bifurcation points and contrasts with the usual exponential evolution of perturbations elsewhere. In our random walk model, however, the relaxation at the edge of intermittency is faster than the one in the “regular” localized phase: the reason is that the latter phase is also critical and characterized by a power-law with a smaller exponent $|\nu|$. Hence it is more adequate to talk here about a critical speed-up when the intermittency threshold is approached.
@@
Many theoretical questions remain pending, such as a proper description of intermittent localization and of the surprisingly simple exponential distribution of localization times. It would be also instructive to investigate the role of memory decay on the properties of the monkey walk in more complex and time dependent environments.
@@ A.1 Power-law kernel
A sample from the probability distribution $\phi(\tau)$ is obtained by equating a uniformly random variable, $u\sim U(0,1)$, to the cumulative distribution function $\int_{0}^{\tau}\phi(\tau^{\prime})d\tau^{\prime}$ and solving for $\tau$. We thus have
@@ Appendix B Phase separation algorithm
For the exponential memory kernel, we implemented an algorithm to decompose a trajectory into localized and diffusive phases. This procedure partitions the trajectories into two interval types, depending on whether the walker occupies often the impurity located at the origin (algorithm 1). Each interval is then classified as localized or de-localized (algorithm 2) based on whether it exceeds a minimum threshold duration, $l_{loc}$ or $l_{deloc}$, respectively. The first $3\Delta$ steps were discarded from all trajectories to eliminate initial transients.
More precisely, given a trajectory of $T$ steps, we determine whether the walker occupies the impurity at the origin by defining a binary variable, $\chi_{t}$, such that $\chi_{t}=1$ when $X_{t}=0$ and $\chi_{t}=0$ otherwise. From the cumulative sum $S_{t}$ of this binary variable, we use the algorithm 1 to identify two types of intervals: increasing segments, corresponding to successive steps with $X_{t}=0$, and constant segments, corresponding to $X_{t}\neq 0$. This algorithm uses the change in the slope of $S_{t}$ between two successive time steps as a criterion to determine the interval type.
@@
For the exponential memory kernel, we define $l_{\text{deloc}}$ in terms of the $\Delta$, the characteristic time scale of the memory range. We arbitrarily classify all constant intervals longer than $l_{deloc}=3\Delta$ as de-localized. In contrast, for the power-law memory kernel, the parameter $\beta$ cannot be directly interpreted as a characteristic scale for the memory. To address this, we compared the curves of the time averaged $P_{0}$ vs. $q$ (up to $t=10^{6}$) between the exponential and power-law memory kernels, identified the ones that were nearly identical and established a linear correspondence between the values of $\Delta$ and $\beta$. Each value for $\beta$ was thus mapped to a corresponding $\Delta$ value, allowing us to define the minimum length of the de-localized intervals as $l_{\text{deloc}}=3\Delta$.
@@
A final filtering step is applied to the localized intervals on the basis of the interval length. Since $1-\gamma$ represents the probability that the walker stays only one unit of time while visiting the impurity, $1/(1-\gamma)$ is the average trapping time. For this reason, we imposed that increasing intervals shorter than $3/(1-\gamma)$ are excluded from the localized phase, as they are not long enough to reflect localization behavior. These intervals are not considered in the analysis of either phase.
@@ Appendix C Relaxation of the perturbation amplitude $h(t)$ for $\beta=1$
We seek to evaluate the second integral in the rhs of Eq. (81). We fix once again a positive $\epsilon\ll 1$ and a large time such that $\epsilon t\gg 1$. Making the change $\tau=t-t^{\prime}$,
@@
In the first contribution AA, we expand the logarithm at first order in $\tau/t$ to obtain its leading behavior

### v3

@@ VIII Conclusions
Previous work has indicated the existence of localized states for certain random walks with preferential resetting to sites visited in the past [20, 21, 22]. A simple model considers an infinite $1d$ lattice, with a single impurity site representing a resource patch with a larger attractiveness than the other sites. Due to the reinforced dynamics, the walker does not diffuse away at late times but tends to an exponential stationary distribution peaked at the impurity. This emergent property can be associated to a phenomenon of spatial learning, mimicking the activity of a foraging animal that would settle down thanks to memory around a zone where resources are abundant in a scarce environment. In the present work, we have studied how such a localization is affected by imperfect memory, when the walker slowly forgets about the steps it performed further in the past. This assumption is more realistic for animals and humans and also gives rise to new phenomena.
@@
As a first result, we have observed that if memory decays slowly enough, such that the effective number of remembered time-steps still grows unbounded with time (albeit sub-linearly), the walker keeps localizing and its stationary density is actually the same as in the perfect memory case. Secondly, according to a de-correlation approximation, these localized solutions turn out to be linearly stable. The relaxation of small perturbations depends on the memory kernel and is not exponential in time: the leading perturbation mode relaxes to zero as an inverse power-law $t^{-|\nu|}$ with a non-trivial exponent which depends on the parameters of the model. We note that a power-law is also thought to describe the relaxation of the classic vertex reinforced random walk, in connection with urn models [7].
@@
Thirdly, a memory loss of the form of $1/\tau$ represents a boundary case which displays interesting peculiarities: it is the memory that decays the fastest and still produce the same localization as the perfect memory case. In other words, this memory is the “cheapest”, since the effective number of remembered steps up to time $t$ only grows as $\ln t$. In addition, somehow counter-intuitively, a walker using the $1/\tau$-kernel exhibits the fastest relaxation toward the localized distribution, see Eq. (61). Hence, such a walker learns faster than a walker with better memory. The speed of learning measured through the exponent $|\nu|$ obtained from Eq. (56) increases as $\beta$, the exponent that describes the decay of memory, varies from 0 to $1$. This result resonates with empirical observations in the neurosciences indicating that forgetting is essential to the brain and can actually be beneficial to learning processes [37, 38, 39]. Memory kernels in $1/\tau$ have been inferred from bison foraging data [32] and can emerge in mean-field models of interacting synapses subject to Hebbian connections and competition [50].
@@
As a fourth result, spatial learning starts to be disrupted when the memory kernel decays exponentially, or faster than $1/\tau$, i.e., when the effective number of remembered steps saturates to a constant value at late times. A new regime of intermittent localization appears, which consists of random time intervals of localization separated by random diffusive excursions that wander away from the impurity. Numerically, the duration of these localized periods is exponentially distributed whereas that of the diffusive excursions follows a power-law distribution with infinite first moment. Consequently, over a very long observation window, the walker will tend to be found mostly in the non-localized state. Nevertheless, as the parameter of memory use ($q$) increases beyond a crossover value, the mean duration of the localized periods becomes extremely large. Remarkably, during a localization period, the distribution of the walker’s position is practically undistinguishable from the localized profile with perfect memory, even when the localization periods are relatively short. Unfortunately, in the intermittent localization regime, the de-correlation approximation breaks down and is not able to predict the loss of stability (or the absence) of stationary solutions.
@@
Intermittency in low-dimensional deterministic dynamical systems is well-known to occur through a saddle-node bifurcation, beyond which an attractor fixed point no longer exists, allowing intermittent bursts of chaos [51]. Just at this bifurcation, the relaxation toward the marginal fixed point follows a power-law form in $1/t$, which is similar although not strictly identical to Eq. (61) for our boundary case. The $1/t$ relaxation at the edge of intermittency in deterministic systems corresponds to a critical slowing down, which is generic at bifurcation points and contrasts with the usual exponential evolution of perturbations elsewhere. In our random walk model, however, the relaxation at the edge of intermittency is faster than the one of the “regular” localized phase: the reason is that the latter phase is also critical and characterized by a power-law with a smaller exponent $|\nu|$. Hence, it is more adequate to talk here about a critical speed-up when the intermittency threshold is approached.
@@
Many theoretical questions remain pending, such as a proper description of intermittent localization and the surprisingly simple exponential distribution of localization times. In addition, it would be instructive to investigate the role of memory decay on the properties of the monkey walk (in particular the critical speed-up effect) in dimension two or larger, or when the underlying process is a Lévy flight. More complex, time dependent environments should also be considered.
@@ A.1 Power-law kernel
A sample from the probability distribution $\phi(\tau)$ is obtained by equating the uniform random variable $u$ in $[0,1]$ to the cumulative distribution function $\int_{0}^{\tau}\phi(\tau^{\prime})d\tau^{\prime}$ and solving for $\tau$. We thus have
@@ Appendix B Phase separation algorithm
For the exponential memory kernel, we implemented an algorithm to decompose a trajectory into localized and diffusive phases. This procedure partitions the trajectories into two interval types, depending on whether the walker often occupies the impurity located at the origin (algorithm 1). Each interval is then classified as localized or de-localized (algorithm 2) based on whether it exceeds a minimum threshold duration, $l_{loc}$ or $l_{deloc}$, respectively. The first $3\Delta$ steps were discarded from all trajectories to eliminate initial transients.
More precisely, given a trajectory of $T$ steps, we determine whether the walker occupies the impurity at the origin by defining a binary variable, $\chi_{t}$, such that $\chi_{t}=1$ when $X_{t}=0$ and $\chi_{t}=0$ otherwise. From the cumulative sum $S_{t}$ of this binary variable, we use the algorithm 1 to identify two types of intervals: increasing segments, corresponding to successive steps with $X_{t}=0$, and constant segments, corresponding to $X_{t}\neq 0$. This algorithm uses the change of the slope of $S_{t}$ between two successive time-steps as a criterion to determine the type of interval.
@@
For the exponential memory kernel, we define $l_{\text{deloc}}$ in terms of the $\Delta$, the characteristic time scale of the memory range. We arbitrarily classify all constant intervals longer than $l_{deloc}=3\Delta$ as de-localized. In contrast, for the power-law memory kernel, the parameter $\beta$ cannot be directly interpreted as a characteristic scale for the memory. To address this, we compared the curves of the time averaged $P_{0}$ vs. $q$ (up to $t=10^{6}$) between the exponential and power-law memory kernels, identified the ones that were nearly superposed and established a linear correspondence between the values of $\Delta$ and $\beta$. Each value of $\beta$ was thus mapped to a corresponding $\Delta$ value, allowing us to define the minimum length of the de-localized intervals as $l_{\text{deloc}}=3\Delta$.
@@
A final filtering step is applied to the localized intervals on the basis of the interval length. Since $1-\gamma$ represents the probability that the walker stays only one unit of time while visiting the impurity, $1/(1-\gamma)$ is the average trapping time. For this reason, we imposed that increasing intervals shorter than $3/(1-\gamma)$ are excluded from the localized phase, as they are not long enough to reflect the localization behavior. These intervals are not considered in the analysis of either phase.
@@ Appendix C Relaxation of the perturbation amplitude $h(t)$ for $\beta=1$
We seek to evaluate the second integral in the rhs of Eq. (84). We fix once again a positive $\epsilon\ll 1$ and a large time $t$ such that $\epsilon t\gg 1$. Making the change $\tau=t-t^{\prime}$,
@@
In the first contribution $A$, we expand the logarithm at first order in $\tau/t$ to obtain its leading behavior